OpenStudy (iwanttogotostanford):

ALGEBRA II HELP: Which of the following is the conjugate of 10 − 3i?

OpenStudy (anonymous):

change the sign in front of the imaginary part to find the conjugate

OpenStudy (anonymous):

(a + bi and (a - bi) are conjugates

OpenStudy (anonymous):

It would be 10+3i For example Conjugates: 3-2i, 3+2i

OpenStudy (campbell_st):

remember part of the general quadratic formula \[-b \pm \sqrt{b^2 - 4ac}\] so if you think about \[3 \pm \sqrt{-4} = 3 \pm 2i\] so the integer doesn't change its just the sign of the imaginary part

OpenStudy (campbell_st):

its also like the difference of 2 squares how do you get a 2 term quadratic (a -b)(a +b) = a^2 - b^2

OpenStudy (iwanttogotostanford):

sorry, my wi-fi went out for awhile @campbell_st @twistnflip @peachpi

OpenStudy (iwanttogotostanford):

so i would want to do the opposite for this one?

OpenStudy (iwanttogotostanford):

what is the first step?

OpenStudy (anonymous):

Yes the conjugates are easy. just switch the sign from positive to negative or vice versa. Here the conjugate is 10+3i

OpenStudy (iwanttogotostanford):

so it would be 13i?

OpenStudy (iwanttogotostanford):

wow I'm being stupid

OpenStudy (anonymous):

Nope do you have answer choices. It wouldn't be 13i because you can't add the two terms because one has a variable and the other does not lol

OpenStudy (anonymous):

The answer is 10 + 3i <-- Final answer

OpenStudy (iwanttogotostanford):

i get it the question was just asking for the conjugate so its 10+3i

OpenStudy (iwanttogotostanford):

my answer choices are

OpenStudy (iwanttogotostanford):

10 + 3i 10 − 3i 3 + 10i 3 − 10i

OpenStudy (anonymous):

a.

OpenStudy (anonymous):

Make sense?

OpenStudy (iwanttogotostanford):

thanks for helping me better understand, everyone!

OpenStudy (iwanttogotostanford):

yep, makes perfect sense